Calculate the energy for the transition of an electron from the n=1 level to the n= 7 level of a hydrogen atom. AE = Joules ICalculate the energy for the transition of an electron from the n=1 level to the n=5 level of a hydrogen atom. AE = Joules IsCalculate the energy of an electron in the n= 3 level of a hydrogen atom. Energy = Joules
Answer #1

From Bohr's model we have equation, E = - 2.178 \times 10 -18 J ( Z 2 / n 2 )

Where, Z is nuclear charge and n is integer.

For Hydrogen atom, Z= 1.

Energy corresponding to state 1 is calculated as

E 1 = - 2.178 \times 10 -18 J ( Z 2 / n 2 ) = - 2.178 \times10 -18 J ( 1 2 / 1 2 ) = - 2.178\times 10 -18 J

Energy corresponding to state 7 is calculated as

E 7 = - 2.178 \times 10 -18 J ( Z 2 / n 2 ) = - 2.178 \times 10 -18 J ( 1 2 / 7 2 ) = - 4.445 \times 10 -20 J

The change in energy when electrons jumps from n=1 to n=7 is

php6hmJRG.png E = energy of final state (E 7 ) - energy of initial state (E 1)

php0GTqAr.pngphp6h9WkB.png E =( - 4.445 \times 10 -20 J ) - ( - 2.178 \times 10 -18 J ) = + 2.134 \times 10 -18 J

+ sign indicates energy is absorbed during electronic transition.

ANSWER : php6hmJRG.png E = + 2.134 \times 10 -18 J

Process : Absorption

PART 2

Energy corresponding to state 1 is calculated as

E 1 = - 2.178 \times 10 -18 J ( Z 2 / n 2 ) = - 2.178 \times 10 -18 J ( 1 2 / 1 2 ) = - 2.178 \times 10 -18 J

Energy corresponding to state 5 is calculated as

E 5 = - 2.178 \times 10 -18 J ( Z 2 / n 2 ) = - 2.178 \times 10 -18 J ( 1 2 / 5 2 ) = - 8.712 \times 10 -20 J

The change in energy when electrons jumps from n=1 to n=5 is

php6hmJRG.png E = energy of final state (E 5 ) - energy of initial state (E 1)

php0GTqAr.pngphp6h9WkB.png E =( - 8.712 \times 10 -20 J ) - ( - 2.178 \times10 -18 J ) = + 2.091 \times 10 -18 J

+ sign indicates energy is absorbed during electronic transition.

ANSWER : php6hmJRG.png E = + 2.091 \times 10 -18 J

Process : Absorption

PART 3

Energy corresponding to state 3 is calculated as

E 3 = - 2.178 \times10 -18 J ( Z 2 / n 2 ) = - 2.178 \times 10 -18 J ( 1 2 / 3 2 ) = - 2.420 \times 10 -19 J

ANSWER : Energy = - 2.420 \times 10 -19 J